OpenStudy (anonymous):

Simplify 10 cubed root of 7 end root minus 7 fourth root of 7 end root minus 9 cubed root of 7 end root plus fourth root of 7

OpenStudy (anonymous):

\[\large 10 \sqrt[3]7 - 7 \sqrt[4]7 -9 \sqrt[3]7 + \sqrt[4]7 \] practicing latexing. is it that?

OpenStudy (anonymous):

your problem i mean...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just like in algebraic expressions we have like terms here... the cube roots are all like terms and the 4th roots are all like terms can you combine them?

OpenStudy (anonymous):

would it be -5 square root of 7?

OpenStudy (anonymous):

no.. your answer will be in terms of cube roots and 4th roots.

OpenStudy (anonymous):

let's try it this way...: let x=cube root 7 let y= fourth root 7 so the problem is like this: 10x - 7y - 9x + y can you simplify that?